# Evaluating and Optimizing Educational Content with Large Language Model Judgments

### v1

@@ Abstract
Creating effective educational materials generally requires expensive and time-consuming studies of student learning outcomes. To overcome this barrier, one idea is to build computational models of student learning and use them to optimize instructional materials. However, it is difficult to model the cognitive processes of learning dynamics. We propose an alternative approach that uses Language Models (LMs) as educational experts to assess the impact of various instructions on learning outcomes. Specifically, we use GPT-3.5 to evaluate the overall effect of instructional materials on different student groups and find that it can replicate well-established educational findings such as the Expertise Reversal Effect and the Variability Effect. This demonstrates the potential of LMs as reliable evaluators of educational content. Building on this insight, we introduce an instruction optimization approach in which one LM generates instructional materials using the judgments of another LM as a reward function. We apply this approach to create math word problem worksheets aimed at maximizing student learning gains. Human teachers’ evaluations of these LM-generated worksheets show a significant alignment between the LM judgments and human teacher preferences. We conclude by discussing potential divergences between human and LM opinions and the resulting pitfalls of automating instructional design.
@@ 1 Introduction
TODO Refer to caption
Building on the insight that LMs can, to some extent, mimic educational experts, along with prior work demonstrating LMs’ capability for iterative improvement [20, 5, 50, 49, 36], we propose an instruction optimization approach (Figure 1). In this approach, one LM (the optimizer) generates new instructional materials, and another LM (the evaluator) evaluates these materials by predicting students’ learning outcomes (e.g., post-test scores). We apply this approach to optimize math word problem worksheets, aiming to maximize students’ post-test scores. In each optimization step, we prompt the optimizer LM to generate new worksheets based on a list of previously generated worksheets with their corresponding post-test scores, then the new worksheets are scored by the evaluator LM. We show that the optimizer LM can refine worksheets based on the judgments of the evaluator LM. We also ask human teachers to compare worksheets generated from different stages of optimization and find a significant correlation between the evaluator LM’s judgments and human preferences. This highlights the potential of LMs in informing the design of real-life experiments and reducing the number of costly experiments in education. However, human teachers sometimes cannot distinguish worksheets that LMs perceive as different, which suggests the necessity for further investigations to ensure LMs effectively complement traditional educational research methodologies.
@@ 3.1 LM-based simulated expert evaluation
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to study a problem and its solution. Here’s the problem: A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? Here’s its solution: Step 1: Identify the amount of sugar needed for the brownie recipe, which is 350 grams. Step 2: Understand that the pound cake recipe requires 270 more grams of sugar than the brownie recipe. Step 3: Add the additional 270 grams of sugar to the 350 grams required for the brownie recipe. Step 4: The total amount of sugar needed for the pound cake recipe is 350 grams + 270 grams = 620 grams. The student is then asked to work on the following problem on their own: The size of a compressed file is 1.74 MiB, while the size of the original uncompressed file is 5.5 times greater. What is the size of the uncompressed file, in MiB? (… more instructions …) Now the student is asked to work on the following problem on a test: Alyssa is twelve years older than her sister, Bethany. The sum of their ages is forty-four. Find Alyssa’s age. Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
@@
TODO Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
@@ 3.3.2 Replication
We create the two types of instructions (practice and worked example), the pre-test, and the post-test, using math word problems from the Algebra dataset [9]. We randomly select 8 problems for the two instructions and a different set of 8 problems $\{x_{1},...,x_{8}\}$ for the pre-test and post-test. The practice instruction (see Figure 15 in Appendix) contains 8 word problems with no solutions, while the worked-example instruction includes the same set of 8 problems with step-by-step solutions for the first, third, fifth, and seventh problems (see Figure 16 in Appendix).
@@ 3.3.3 Results
TODO Refer to caption
@@ 3.4.2 Replication
We create four types of instructions (one for each experimental group) and a post-test, using math word problems from the Algebra dataset [9]. We randomly select a set of 6 problems $\{x_{1},...,x_{6}\}$ for the post-test, which is identical in all conditions. In the instructions for both low-variability and high-variability conditions, the first, third, and fifth problems are identical. In the low-variability conditions, the second, fourth, and sixth problems have the same format as the odd-numbered problems but use different values. In the high-variability conditions, the even-numbered problems differ from the odd-numbered ones in both values and formats. In practice conditions, only problems are presented. In worked-example conditions, all problems and step-by-step solutions are presented (see Appendix B.3). We simulate post-tests as described in Section 3.2. We use gpt-3.5-turbo-16k as the evaluator LM for all SEEs, with the temperature parameter set to 0 for more consistent output.
@@ 3.4.3 Results
TODO Refer to caption
@@ 4.1 The Instruction Optimization algorithm
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5, where higher numbers indicate more proficiency): 1. Being able to set up systems of equations given a word problem: 1 2. Being able to solve systems of equations: 1 I have some worksheets along with the student’s test scores after receiving the worksheets. The worksheets are arranged in ascending order based on their scores, where higher scores indicate better quality. Worksheet: You need to study a problem and its solution. Here’s the problem: A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? Here’s its solution: Step 1: Identify the amount of sugar needed for the brownie recipe, which is 350 grams. Step 2: Understand that the pound cake recipe requires 270 more grams of sugar than the brownie recipe. Step 3: Add the additional 270 grams of sugar to the 350 grams required for the brownie recipe. Step 4: The total amount of sugar needed for the pound cake recipe is 350 grams + 270 grams = 620 grams. Test score: 20 Worksheet: {worksheet2} Test score: 34 Worksheet: {worksheet3} Test score: 66 (…more worksheets and scores…) Generate a new worksheet to further increase the test score of the student. You will be evaluated based on this score function: “‘python {utility_string} ”’ The new worksheet should begin with <WORKSHEET> and end with </WORKSHEET>.
@@ 4.2 Optimizing math worksheets
TODO Refer to caption
@@ 4.3 Human evaluation of worksheets
We asked participants to do pairwise comparisons where they need to indicate their preferences between two worksheets (worksheet A and B) using a continuous slider scale, with a rating of $-1$ assigned for a strong preference towards the first worksheet, a rating of $1$ for a strong preference towards the second worksheet, and a rating of 00 for no preference (see the task description in Figure 22 in Appendix).
@@
TODO Refer to caption
Figure 10: Evaluations by human teachers show a significant correlation between the LM judgments and human teacher preferences (Pearson $r=0.661,p<0.05$). The error bars are bootstrapped 95% confidence intervals.
5 discussion and conclusion
We show that LMs can act as evaluators on a couple of well-known findings for pedagogical effectiveness, including results that require understanding that different students can be impacted differently by different content. Even if the LM succeeds at these judgments due to reading relevant studies, it is still remarkable for two reasons: 1) Nothing in the evaluations directly cued those studies, and 2) We developed new instructional materials in a domain that is different from the ones explored in prior studies. This shows that the LM can read and synthesize educational science reliably into useful evaluators. In addition, LM evaluators correlated positively with expert humans in their evaluations of the effectiveness of new content.
However, there may be other known educational findings that LM evaluators do not replicate, and human teachers sometimes cannot distinguish worksheets that LMs perceive as different and may disagree with LMs on what is optimal. Therefore, the LM may be useful as a coarse evaluator to help speed up content design as a human augmentor but not to replace human evaluators or empirical studies. An interesting open question is whether we can extend LM-based simulated educational experts to multi-modal instructional input.
@@ Appendix A Modeling cognitive processes of learning dynamics
TODO You are a student who is initially ignorant. You are learning by watching a video. Here’s the video: In this video, we’re gonna get some more practice setting up systems of equations. So we’re told Sanjay’s dog weighs five times as much as his cat. His dog is also 20 kilograms heavier than his cat. Let’s stop the video. Remember, you’ve only been taught what was shown in the video. It typically takes you 5-10 practices to learn a new skill, and you often need to rewatch a lecture and do practice problems. Remember you have done 0 examples, and that is less than 5! Can you apply what you’ve just learned to solve the following problem? If you do not know how to solve it, don’t try to guess and ask to keep watching the video. If you do know how to solve it, then place the final answer (a number) in square brackets. Here’s the problem: Alyssa is twelve years older than her sister, Bethany. The sum of their ages is forty-four. Find Alyssa’s age.
@@
TODO Refer to caption
Figure 13: The first 30 sentences in the video transcript.
@@ Appendix B Prompts
B.1 Pre-test prompt
@@
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
@@ B.2 Replicating the Expertise Reversal Effect
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to work on the following problems on their own: 1. A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? 2. The size of a compressed file is 1.74 MiB, while the size of the original uncompressed file is 5.5 times greater. What is the size of the uncompressed file, in MiB? 3. Devon is 26 years older than his son Cooper. The sum of their ages is 50. Find Devon’s age. 4. When Jenna spent 10 minutes on the elliptical trainer and then did circuit training for 20 minutes, her fitness app says she burned 278 calories. When she spent 20 minutes on the elliptical trainer and 30 minutes circuit training she burned 473 calories. How many calories does she burn for each minute on the elliptical trainer? 5. Last month Jim and Debbie earned \$7,200. Debbie earned \$1,600 more than Jim earned. How much did Debbie earn? 6. Pam is 3 years older than her sister, Jan. The sum of their ages is 99. Find Pam’s age. 7. Lynn paid a total of \$2,780 for 261 tickets to the theater. Student tickets cost \$10 and adult tickets cost \$15. How many student tickets did Lynn buy? 8. Yumi wants to make 12 cups of party mix using candies and nuts. Her budget requires the party mix to cost her \$1.29 per cup. The candies are \$2.49 per cup and the nuts are \$0.69 per cup. How many cups of candies should she use? Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
Figure 15: A prompt for the practice group in simulated expert evaluations replicating the Expertise Reversal Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM.
@@
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to study a problem and its solution. Here’s the problem: A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? Here’s its solution: Step 1: Identify the amount of sugar needed for the brownie recipe, which is 350 grams. Step 2: Understand that the pound cake recipe requires 270 more grams of sugar than the brownie recipe. Step 3: Add the additional 270 grams of sugar to the 350 grams required for the brownie recipe. Step 4: The total amount of sugar needed for the pound cake recipe is 350 grams + 270 grams = 620 grams. The student is then asked to work on the following problem on their own: The size of a compressed file is 1.74 MiB, while the size of the original uncompressed file is 5.5 times greater. What is the size of the uncompressed file, in MiB? The student is then asked to study another problem and its solution. Here’s the problem: Devon is 26 years older than his son Cooper. The sum of their ages is 50. Find Devon’s age. Here’s its solution: Step 1: Let’s denote Devon’s age as D and Cooper’s age as C. Step 2: According to the problem, we have two equations. The first one is D = C + 26 (since Devon is 26 years older than Cooper). The second one is D + C = 50 (since the sum of their ages is 50). Step 3: We can substitute the first equation into the second one. So, instead of D in the second equation, we can put C + 26. We get C + 26 + C = 50. Step 4: Simplify the equation. Combine like terms to get 2C + 26 = 50. Step 5: Subtract 26 from both sides of the equation to get 2C = 24. Step 6: Divide both sides of the equation by 2 to get C = 12. So, Cooper is 12 years old. Step 7: Substitute C = 12 into the first equation (D = C + 26) to get D = 12 + 26 = 38. So, Devon is 38 years old. The student is then asked to work on the following problem on their own: When Jenna spent 10 minutes on the elliptical trainer and then did circuit training for 20 minutes, her fitness app says she burned 278 calories. When she spent 20 minutes on the elliptical trainer and 30 minutes circuit training she burned 473 calories. How many calories does she burn for each minute on the elliptical trainer? (… more problems and solutions …) Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
Figure 16: A prompt for the worked-example group in simulated expert evaluations replicating the Expertise Reversal Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM. Note that the first and third problems have step-by-step solutions, while the second and fourth problems do not.
B.3 Replicating the Variability Effect
B.3.1 Low-variability/practice group
[…]
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to work on the following problems on their own: 1. A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? 2. A cookie recipe is asking for 400 grams of sugar, and a brownie recipe requires 70 more grams of sugar than a cookie recipe. How many grams of sugar are needed for the brownie? 3. Devon is 26 years older than his son Cooper. The sum of their ages is 50. Find Devon’s age. 4. Pam is 3 years older than her sister, Jan. The sum of their ages is 99. Find Pam’s age. 5. Last month Jim and Debbie earned \$7,200. Debbie earned \$1,600 more than Jim earned. How much did Debbie earn? 6. A married couple together earn \$75,000. The husband earns \$15,000 more than five times what his wife earns. What does the wife earn? Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
Figure 17: A prompt for the low-variability/practice group in simulated expert evaluations replicating the Variability Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM. Note that problems with even numbers have the same format as problems with odd numbers.
B.3.2 Low-variability/worked example group
[…]
Figure 18: A prompt for the low-variability/worked-example group in simulated expert evaluations replicating the Variability Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM. Note that problems with even numbers have the same format as problems with odd numbers. All problems are presented with their step-by-step solutions.
B.3.3 High-variability/practice group
[…]
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to work on the following problems on their own: 1. A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? 2. The size of a compressed file is 1.74 MiB, while the size of the original uncompressed file is 5.5 times greater. What is the size of the uncompressed file, in MiB? 3. Devon is 26 years older than his son Cooper. The sum of their ages is 50. Find Devon’s age. 4. When Jenna spent 10 minutes on the elliptical trainer and then did circuit training for 20 minutes, her fitness app says she burned 278 calories. When she spent 20 minutes on the elliptical trainer and 30 minutes circuit training she burned 473 calories. How many calories does she burn for each minute on the elliptical trainer? 5. Last month Jim and Debbie earned \$7,200. Debbie earned \$1,600 more than Jim earned. How much did Debbie earn? 6. Lynn paid a total of \$2,780 for 261 tickets to the theater. Student tickets cost \$10 and adult tickets cost \$15. How many student tickets did Lynn buy? Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
Figure 19: A prompt for the high-variability/practice group group in simulated expert evaluations replicating the Variability Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM. Note that problems with even numbers are different than problems with odd numbers in both formats and values.
B.3.4 High-variability/worked example group
[…]
TODO Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5): 1. Being able to set up systems of equations given a word problem: {level1} 2. Being able to solve systems of equations: {level2} Here’s the instruction that the student receives. The student is asked to study a problem and its solution. Problem 1: A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? Solution: 1. Identify the amount of sugar needed for the brownie recipe, which is 350 grams. 2. Understand that the pound cake recipe requires 270 more grams of sugar than the brownie recipe. 3. Add the additional 270 grams of sugar to the 350 grams required for the brownie recipe. 4. The total amount of sugar needed for the pound cake recipe is $350+270=620$ grams. The student is then asked to study another problem and its solution. Here’s the problem: Problem 2: The size of a compressed file is 1.74 MiB, while the size of the original uncompressed file is 5.5 times greater. What is the size of the uncompressed file, in MiB? Solution: 1. Identify the size of the compressed file, which is 1.74 MiB. 2. Identify the ratio of the uncompressed file size to the compressed file size, which is 5.5. 3. Multiply the size of the compressed file by the ratio to find the size of the uncompressed file. Thus, $1.74\times 5.5=9.57$ MiB. (… more problems and solutions …) Now the student is asked to work on the following problem on a test: {problem} Given the student’s initial skill levels and the instruction the student has received, what’s the probability that the student can solve the problem correctly? Explain your reasoning and give a single number between 0 and 100 in square brackets.
Figure 20: A prompt for the high-variability/worked-example group in simulated expert evaluations replicating the Variability Effect. The blue text is the student persona. The black text is the instruction given to the student. The orange text is a problem on the post-test. The green text describes the evaluation task for the LM. Note that problems with even numbers are different than problems with odd numbers in both formats and values. All problems are presented with their step-by-step solutions.
B.4 Optimizing math word problem worksheets
[…]
TODO Refer to caption
Appendix C Human teacher evaluations
[…]
TODO Refer to caption
Figure 22: The evaluation task description for teachers. Teachers can indicate their preference by adjusting the slider.
[…]

### v2

@@ Abstract
Creating effective educational materials generally requires expensive and time-consuming studies of student learning outcomes. To overcome this barrier, one idea is to build computational models of student learning and use them to optimize instructional materials. However, it is difficult to model the cognitive processes of learning dynamics. We propose an alternative approach that uses Language Models (LMs) as educational experts to assess the impact of various instructions on learning outcomes. Specifically, we use GPT-3.5 to evaluate the overall effect of instructional materials on different student groups and find that it can replicate well-established educational findings such as the Expertise Reversal Effect and the Variability Effect. This demonstrates the potential of LMs as reliable evaluators of educational content. Building on this insight, we introduce an instruction optimization approach in which one LM generates instructional materials using the judgments of another LM as a reward function. We apply this approach to create math word problem worksheets aimed at maximizing student learning gains. Human teachers’ evaluations of these LM-generated worksheets show a significant alignment between the LM judgments and human teacher preferences. We conclude by discussing potential divergences between human and LM opinions and the resulting pitfalls of automating instructional design.  [^1] Our prompts are publicly available at https://github.com/StanfordAI4HI/ed-expert-simulator.
@@ 1 Introduction
We use an optimizer LM to generate new educational materials (e.g., worksheets) and an evaluator LM to judge the effectiveness of these materials by predicting students’ post-test performance. In this example, the optimizer LM generates a new worksheet based on the student information and previous worksheet-score pairs, then the evaluator LM predicts the student’s post-test score given the new worksheet. The new worksheet-score pair serves as feedback for the optimizer LM to refine worksheets.
Building on the insight that LMs can, to some extent, mimic educational experts, along with prior work demonstrating LMs’ capability for iterative improvement [20, 5, 50, 49, 36], we propose an instruction optimization approach (Figure 1). In this approach, one LM (the optimizer) generates new instructional materials, and another LM (the evaluator) evaluates these materials by predicting students’ learning outcomes (e.g., post-test scores). We apply this approach to optimize math word problem worksheets, aiming to maximize students’ post-test scores. External assessments like post-tests are not perfect, but they are an important tool in many educational systems to help shed insight into student progress. In each optimization step, we prompt the optimizer LM to generate new worksheets based on a list of previously generated worksheets with their corresponding post-test scores, then the new worksheets are scored by the evaluator LM. We show that the optimizer LM can refine worksheets based on the judgments of the evaluator LM. We also ask human teachers to compare worksheets generated from different stages of optimization and find a significant correlation between the evaluator LM’s judgments and human preferences. This highlights the potential of LMs in informing the design of real-life experiments and reducing the number of costly experiments in education. However, human teachers sometimes cannot distinguish worksheets that LMs perceive as different, which suggests the necessity for further investigations to ensure LMs effectively complement traditional educational research methodologies.
@@ 3.3.2 Replication
We create the two types of instructions (practice and worked example), the pre-test, and the post-test, using math word problems from the Algebra dataset [9]. We randomly select 8 problems for the two instructions and a different set of 8 problems $\{x_{1},...,x_{8}\}$ for the pre-test and post-test. The practice instruction contains 8 word problems with no solutions, while the worked-example instruction includes the same set of 8 problems with step-by-step solutions for the first, third, fifth, and seventh problems. Detailed instructions are available at https://github.com/StanfordAI4HI/ed-expert-simulator.
@@ 3.3.3 Results
The figure presents two line graphs comparing the scores of low-expertise and high-expertise learners before and after an intervention, which is either a practice or worked-example instruction.
@@ 3.4.2 Replication
We create four types of instructions (one for each experimental group) and a post-test, using math word problems from the Algebra dataset [9]. We randomly select a set of 6 problems $\{x_{1},...,x_{6}\}$ for the post-test, which is identical in all conditions. In the instructions for both low-variability and high-variability conditions, the first, third, and fifth problems are identical. In the low-variability conditions, the second, fourth, and sixth problems have the same format as the odd-numbered problems but use different values. In the high-variability conditions, the even-numbered problems differ from the odd-numbered ones in both values and formats. In practice conditions, only problems are presented. In worked-example conditions, all problems and step-by-step solutions are presented. We use gpt-3.5-turbo-16k as the evaluator LM for all SEEs, with the temperature parameter set to 0 for more consistent output.
@@ 3.4.3 Results
The vertical axis represents post-test scores, ranging from 55.0 to 75.0.
@@
[…]
The vertical axis represents post-test scores, ranging from 40.0 to 80.0.
@@ 4.1 The Instruction Optimization algorithm
Here is an 8th-grade student with the following skill levels (each skill is rated on a scale from 1 to 5, where higher numbers indicate more proficiency): 1. Being able to set up systems of equations given a word problem: 1 2. Being able to solve systems of equations: 1 I have some worksheets along with the student’s test scores after receiving the worksheets. The worksheets are arranged in ascending order based on their scores, where higher scores indicate better quality. Worksheet: You need to study a problem and its solution. Here’s the problem: A brownie recipe is asking for 350 grams of sugar, and a pound cake recipe requires 270 more grams of sugar than a brownie recipe. How many grams of sugar are needed for the pound cake? Here’s its solution: Step 1: Identify the amount of sugar needed for the brownie recipe, which is 350 grams. Step 2: Understand that the pound cake recipe requires 270 more grams of sugar than the brownie recipe. Step 3: Add the additional 270 grams of sugar to the 350 grams required for the brownie recipe. Step 4: The total amount of sugar needed for the pound cake recipe is 350 grams + 270 grams = 620 grams. Test score: 20 Worksheet: {worksheet2} Test score: 34 Worksheet: {worksheet3} Test score: 66 (…more worksheets and scores…) Generate a new worksheet to further increase the test score of the student. You will be evaluated based on this score function: “‘python {utility_string} ”’ The new worksheet should begin with <WORKSHEET> and end with </WORKSHEET>.
@@
The optimization prompt for generating a new instruction (e.g., a worksheet) based on prior instructions and post-test scores for a given student. The blue text is the student persona. The orange text contains the previous worksheet-score pairs, sorted in ascending order. The green text describes the optimization task for the LM. The score function is a Python program that illustrates how post-test scores are computed without revealing the content of the post-test.
@@ 4.2 Optimizing math worksheets
This plot shows a line graph depicting the relationship between the number of optimization steps and the LM-predicted student post-test scores. The horizontal axis of the graph represents the number of optimization steps, which range from 0 to 18. The vertical axis measures the LM-predicted student post-test scores, ranging from 20 to 90. The line graph starts at a lower score at zero optimization steps and generally trends upward as the number of optimization steps increases.
@@ 4.3 Human evaluation of worksheets
We asked participants to do pairwise comparisons where they need to indicate their preferences between two worksheets (worksheet A and B) using a continuous slider scale, with a rating of $-1$ assigned for a strong preference towards the first worksheet, a rating of $1$ for a strong preference towards the second worksheet, and a rating of 00 for no preference (see the task description in Figure 10).
Figure 10: The evaluation task description for teachers. In each evaluation, we ask the teacher to compare two worksheets in terms of their effectiveness in improving the student’s test performance, and we show the teacher two sample test questions. Teachers can indicate their preference by adjusting the slider.
[…]
This outlines a human evaluation task related to assessing the utility of math worksheets for an 8th-grade student.
@@
Figure 11: Evaluations by human teachers ($n=56$) show a significant correlation between the LM judgments and human teacher preferences (Pearson $r=0.661,p<0.05$). The error bars are bootstrapped 95% confidence intervals.
@@
This is a scatter plot comparing LM-predicted student post-test scores with human preference scores.
5 Limitations
There are several limitations to our work. First, the LMs we used were likely trained on data containing descriptions of experiments similar to those we replicated. Therefore, we developed instructional materials in a domain different from previous studies.
Second, one potential concern is that LMs do not consistently replicate established educational results. Our experiments focused solely on mathematics and only tested the performance of GPT-3.5 and GPT-4 with a limited number of straightforward prompts. Exploring all available LMs would have been prohibitively costly, so we chose GPT-3.5 and GPT-4 due to their popularity and robust performance. We used GPT-3.5 as the evaluator LM because it is more cost-effective for handling long prompts. GPT-4 served as our optimizer LM, chosen for its capability in critical assessment, a feature typically reliable only in the most advanced models. Our goal was to demonstrate the potential of LMs in these tasks, not to claim these LM choices as the definitive best.
Third, there are discrepancies between the evaluations made by LMs and those by human teachers, who sometimes fail to perceive differences that LMs identify. These discrepancies raise concerns about the models’ decision-making processes and the potential impact of biases in their training data. Furthermore, our evaluations are limited to the perspective of teachers. However, ultimately, the effectiveness of the educational materials depends on how students engage with them. While LMs could aid in accelerating the design process by providing preliminary evaluations, they should not be seen as replacements for direct evaluations involving human students.
6 Conclusion
We show that LMs can act as evaluators on a couple of well-known findings for pedagogical effectiveness, including results that require understanding that different students can be impacted differently by different content. Even if the LM succeeds at these judgments due to reading relevant studies, it is still remarkable for two reasons: 1) Nothing in the evaluations directly cued those studies, and 2) We developed new instructional materials in a domain different from the ones explored in prior studies. This shows that the LM can read and synthesize educational science reliably into useful evaluators. In addition, LM evaluators correlated positively with expert humans in their evaluations of the effectiveness of new content.
However, there may be other known educational findings that LM evaluators do not replicate, and human teachers sometimes cannot distinguish instructional materials that LMs perceive as different and may disagree with LMs on what is optimal. Therefore, the LM may be useful as a coarse evaluator to help speed up content design as a human augmentor but not to replace human evaluators or empirical studies. An interesting open question is whether we can extend LM-based simulated educational experts to multi-modal instructional input.
7 Acknowledgements
We would like to thank Dongwei Jiang, Rose E. Wang, Allen Nie, and Shubhra Mishra for their feedback on our work. This work was supported by the Junglee Corporation Stanford Graduate Fellowship and a Stanford Hoffman-Yee grant.
@@ Appendix A Modeling cognitive processes of learning dynamics
shows a comparison of accuracy rates between two models, GPT-4-0613 and GPT-3.5 Turbo, across a range of sentence counts. The horizontal axis of the graph indicates the number of sentences, ranging from 0 to 40. The vertical axis measures accuracy, ranging from 0.0 to 0.8.
@@ Appendix B Prompts
Prompts used in this work are publicly available at https://github.com/StanfordAI4HI/ed-expert-simulator.
@@
see caption.